Sample Title

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2003 Jorgen Skogstad

1 First Chapter

This is the content of the first chapter. Here is a numbered list.

1. This is the first item

2. This is the second item The makeinfo and texinfo-format-buffer commands transform a Texinfo file such as

this into a Info file; and TEX typesets it for a printed manual.

2 Texinfo template Concept Index 3

Concept Index

A

Another sample index entry . . . . . . . . . . . . . . . . . . . 1

S

Sample index entry . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

4 Texinfo template

Table of Contents

1 First Chapter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

ii Texinfo template